OpenStudy (anonymous):

find the sum of the series (1/2^k + 1/4^k)

OpenStudy (psymon):

The addition of geometric series. I figure you can just take the sum of each series individually and then add them? Either way, since these are geo-series they fit the form of: \[\sum_{i=0}^{\infty}ar^{n}\]A would be your starting point so to speak and r would be your ratio. In this case, your a is just 1 really and the r is 1/2. So you have: \[1(\frac{ 1 }{ 2 })^k\]When you have something like this, the sum of the series is \[\frac{ a }{ 1-r }\]So that being said, the first series just gives you: \[\frac{ 1 }{ 1-\frac{ 1 }{ 2 } }= 2\]This then gets added to the other series which would just be: \[\frac{ 1 }{ 1-\frac{ 1 }{ 4 } }= \frac{ 4 }{ 3 }\]So then its just the addition of these 2 \[2+\frac{ 4 }{ 3 }= \frac{ 10 }{ 3 }\]